OpenStudy (anonymous):

17. Model the function rule y = -2x - 3 with a table of values. 18. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

jimthompson5910 (jim_thompson5910):

For 17, you'll be plugging in various values of x to get corresponding values of y if x = 0, then y = -2x - 3 y = -2(0) - 3 ... replace x with 0, then evaluate y = 0 - 3 y = -3 So we'll have this row 

jimthompson5910 (jim_thompson5910):

if x = 1, then y = -2x - 3 y = -2(1) - 3 ... now replace x with 1, since x = 1 y = -2 - 3 y = -5 another pair is x = 1, y = -5